OpenStudy (mathisfun13):

Translate the equation: 19 less than five times a number is equal to 11 more than three times the number

OpenStudy (mathisfun13):

i think its 5x-19=11(x+3)?

OpenStudy (petiteme):

i think its 5x-19=11+3x